OpenStudy (anonymous):

If the tangent line to the graph of the funtion f(x) at the point (1,7) passes through the point (-2,2), then what is f'(x)?

OpenStudy (anonymous):

find the slope of that line i assume it means what is \[f'(1)\]

OpenStudy (anonymous):

you cannot know \[f'(x)\] unless you know \[f(x)\] but \[f'(1)\] will be your slope

OpenStudy (anonymous):

so is it 5/3?

OpenStudy (anonymous):

yes again that is if you were asked for \[f'(1)\]

OpenStudy (anonymous):

oh shoot, i thought i put f'(1). hahaha. thanks satellite73!

OpenStudy (anonymous):

i got the answer from just getting the slope of the two points. i need to show work though. can you help me?

OpenStudy (anonymous):

show the work for finding the slope?

OpenStudy (anonymous):

you can do it with your eyeballs, over 3 up 5 slope is \[\frac{5}{3}\] like you already said. that is all the work. however you computed the slope, write that down and you are done

OpenStudy (anonymous):

something like \[f'(1)=\frac{7-2}{1+2}=\frac{5}{3}\]

OpenStudy (anonymous):

hahaha, i just expected the question to have lots of work to be shown. thanks again!